Please explain these two questions for me,why the option is right and why the others are false. I need your detailed explanation, please type it down. I will thank you forever.

13-6) If the LM curve is vertical, then A) there is partial crowding out of an increase in government expenditures. B) the
Answer #1

1) Option A is not correct because if LM curve is vertical then in IS LM model fiscal policy become totally ineffective so , there is full crowding out (not partial crowding out) . Option B is not correct because if money supply will increase LM curve will shift to the right and will increase GDP with lower interest rate . Option C is not correct because the LM curve is derived from the equilibrium in money market , if money demand does not depend on the interest rate (that is insensitive to change in interest rate ) then LM equation can be written as M/P= L(Y) , which clearly shows LM curve is also independent of interest rate .

Option D is correct because if you look at Quantity theory of money i.e. M*V =P*Y , to have one to one relation between M/P and Y , velocity of money need to remain constant

2) For the money creation process Bank need to have continuous depositing and giving out loan process .

option A and B are not correct because if individual (whether he is a borrower or depositor) keeps reserves with themselves then that part of money cant be used for money creation . So bank always prefer to accepts deposits and keep reserves and lend rest of the money to the borrower (not in cash , it first creates borrowers new account , deposit that borrowed amount and then again keep reserve and lend rest of the money ....money creation process) All this is possible only if paper deposits are acceptable in the market as a mode of payment . That is why option D is correct